OpenStudy (anonymous):

B. If this experiment is repeated at a place 10 kilometers above the sea-level, the acceleration due to gravity g (exp) is expected to: a) increase. b) decrease. c) remain constant.

OpenStudy (akashdeepdeb):

Which experiment? Acceleration is always constant for any height or depth (h << 6400) When h has an appreciable magnitude, it always decreases, g is never greater then 9.8 at a height or at a depth.

OpenStudy (anonymous):

Freely falling bodies experiment ... So acceleration is constant right ???

OpenStudy (mrnood):

the full equation for gravitation force on an object close to earth is F - GMm/r^2 if you take r = radius of earth (i.e. the body is on the earth's surface then we normally simplify this to F = mg (so g= GM/r^2) Now - the radius of the earth is approx. 6.4x10^6 m so for all experiments close to the earth we normally consider g to be essentially constant. in your question you increase r to r+10^4m so you can see that the force of gravity will decrease (1/r^2) However the increase in radius is still very small compared to the radius - so the effect will be very small.

OpenStudy (anonymous):

Thanks for both of you!!!!! :)